OpenStudy (anonymous):

5x-3 all over 11 = -3

OpenStudy (imtant):

Bring 11 to the right and times with -3 you get,5x-3=-33 so next,bring -3 to the right become 5x=-33+3 solve it,5x=-30 and x=-6.

OpenStudy (anonymous):

thankyou :) it means alot lol

OpenStudy (imtant):

Yw :)